Circular No. 8080

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

XTE J1720–318 R. González-Riestra, P. M. Rodríguez-Pascual, M. Santos-Lleó, N.

Schartel, P. Calderón, P. Munuera, and B. Juárez, XMM-Newton Science Operations Centre, report that an XMM-Newton target-of-opportunity observation of XTE J1720–318 (IAUC 8050) on Feb. 20.4 UT shows a bright source at α = 17^{h}19^{m}59^{s}.0, δ = −31^{o}44′59′′ (equinox 2000.0; uncertainty ∼ 4′′) during a 17000-s exposure. This position is consistent with the coordinates of the radio (IAUC 8054) and infrared counterparts (IAUC 8056). The rates were 14 counts/s in the RGS first-order spectrum, 588 counts/s in EPIC-pn (core piled-up), and 346 counts/s in EPIC-MOS1 (timing mode). The spectrum is consistent with a hydrogen column density on the order of 1.3 × 10^{22} cm^{−2}. The observed flux in the range 2–10 keV is 1.6 × 10^{−9} erg cm^{−2} s^{−1}.

XTE J1807–294 C. B. Markwardt, University of Maryland and Goddard Space Flight

Center (GSFC); and E. Smith and J. H. Swank, GSFC, report that RXTE PCA monitoring observations of the Galactic-center region have revealed a new millisecond pulsar, designated XTE J1807–294. PCA cross scans on Feb. 21.6 UT determined the position to be α = 18^{h}07^{m}.0, δ = −29^{o}24′ (equinox 2000.0), with an uncertainty of ≈ 1′ due to systematic errors. Coherent pulsations were detected at a barycentered frequency of 190.625 Hz. While the pointing is interrupted by scans, there appears to be a significant sinusoidal modulation of the pulse frequency, at a period of 35 ± 3 min, and with a half-amplitude of 2.2 ± 0.6 mHz. This source is the fourth known accreting millisecond pulsar, and the orbital period, if confirmed, is the shortest yet of that group. The 2–10-keV x-ray fluxes (in mCrab) were: Feb. 16.7, 33; 19.8, 38; 21.6, 58; 22.6, 41; 25.4, 32 — suggesting that the x-ray peak has passed. Further RXTE observations are planned. Observations at other wavelengths are encouraged.

CORRIGENDA On IAUC 7961, ‘Comet P/2002 Q4 (Brewington)’, first line, for G.

Trancredi, read G. Tancredi,

On IAUC 8079, ‘Comet C/2002 Y1 (Juels-Holvorcem)’, last line, for 10×50 binoculars). read 20×100 binoculars).

2003 February 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green